Write a quadratic equation and solve. Use a calculator and the decimal approximation of 3.14 for π(pi).
The volume of the circular cylinder is 45,600 ft^3. Find its radius to the nearest foot. (the cylinder is 30 ft)
FORMULA:
V = πr^2H

Answer by Cromlix(4381)

Hi there,
Volume = π x r^2 x h
45,600 = π x r^2 x 30
r^2 = 45,600/(π30
r^2 = 483.831027
r = √483.831027
r = 21.996 ft
Radius = 22 ft. (Nearest foot)
Hope this helps :-)